# An Eulerian Vortex Method on Flow Maps

### v1

@@ Abstract.
We present an Eulerian vortex method based on the theory of flow maps to simulate the complex vortical motions of incompressible fluids. Central to our method is the novel incorporation of the flow-map transport equations for line elements, which, in combination with a bi-directional marching scheme for flow maps, enables the high-fidelity Eulerian advection of vorticity variables. The fundamental motivation is that, compared to impulse $\bm{m}$, which has been recently bridged with flow maps to encouraging results, vorticity $\bm{\omega}$ promises to be preferable for its numerically stability and physical interpretability. To realize the full potential of this novel formulation, we develop a new Poisson solving scheme for vorticity-to-velocity reconstruction that is both efficient and able to accurately handle the coupling near solid boundaries. We demonstrate the efficacy of our approach with a range of vortex simulation examples, including leapfrog vortices, vortex collisions, cavity flow, and the formation of complex vortical structures due to solid-fluid interactions.
@@
Figure 1. Turbulent flow generated from a delta wing with 0 viscosity and an angle-of-attack of $20^{\circ}$. Two pictures on the right illustrate smoke (top) and vorticity (bottom) snapshots at different time steps. More stable spiral vortex structure can be generated with viscosity, as shown in Figure 17.
@@ 1. Introduction
Vortex methods have established themselves as a foundational computational method for incompressible flows with turbulence and complex vortical structures. In these methods, the Navier-Stokes equations written in velocity $\bm{u}$ are reformulated into their vorticity form, establishing vorticity
@@ 2. Related Work
Vortex methods rewrite the incompressible Navier-Stokes equation by treating vorticity as a gauge variable for velocity. By explicitly advecting the vorticity, this approach naturally preserves fluid circulation. In order to reduce the numerical dissipation during the advection, diverse vorticity representations have been proposed, ranging from particles (Selle et al., 2005; Cottet et al., 2000; Park and Kim, 2005; Zhang and Bridson, 2014; Angelidis, 2017), to filaments (Angelidis and Neyret, 2005; Weißmann and Pinkall, 2010; Ishida et al., 2022; Padilla et al., 2019), segments (Xiong et al., 2021), sheets (Brochu et al., 2012; Pfaff et al., 2012; Da et al., 2015), and Clebsch level sets (Chern et al., 2016, 2017; Yang et al., 2021; Xiong et al., 2022). For these Lagaragian vorticity representations, the velocity reconstruction from vorticity is often done with the Biot-Savart law, accelerated with the Fast Multipole Method (FMM) (Greengard and Rokhlin, 1987). Despite the enhanced ability to preserve the vortex structure, these vorticity representations are not as convenient to implement as the Eulerian vortex method when handling solid boundaries. Vortex methods with the vorticity-streamfunction formulation have been adopted by several researchers (Zhang et al., 2015; Elcott et al., 2007; Ando et al., 2015; Yin et al., 2023a). In spite of the ability to handle solid boundaries, these methods require solving a potential component of the velocity. Existing Eulerian vortex methods that use the velocity-vorticity formulation avoid the streamfunction but suffer from poor vorticity preservation and inaccurate solid boundary conditions. For example, Huang and Li (1997) only addressed 2D cases and imposed vorticity boundary conditions, rather than velocity. Similarly, Liu (2001) set the vorticity boundary condition simply as $\bm{\omega}_{b}=0$. Our method bypasses the use of streamfunction and potential without sacrificing the correctness in solid boundary handling, and pushes forward the state-of-the-art in vortex method simulations through the novel combination with flow map advection.
@@ 3.1. Flow map
Here, Eqn. (6) defines the forward evolution, and Eqn. (7) defines the time-reversed evolution. In Eqn (6), we have $\bm{\phi}\left(\bm{x}\right)$ and $\bm{x}$ specifying the start and end point of the flow map path at time 00 and $t$. Conversely, in Eqn (7), by examining the backward path from time $t$ to time 00, we can construct a mapping with $\bm{l}\left(\bm{X},t\right)$ denoting the line element located in $\bm{X}$ at time 00 and $\bm{l}\left(\bm{\phi}\left(\bm{X}\right),0\right)$ specifying its value at time $t$.
@@
(10) $\bm{s}(\bm{X},0)=\mathcal{F}^{T}\,\bm{s}\left(\bm{\phi}\left(\bm{X}\right),0\right).$
@@ 3.3. Vorticity on Flow Maps
Here, in Eqn. (15), $\bm{\omega}\left(\bm{x},t\right)$ denotes the vorticity of a material point located in $\bm{x}$ in time $t$ and $\bm{\omega}\left(\bm{\psi}\left(\bm{x}\right),0\right)$ is the vorticity of the same material point in time 00. Conversely, in Eqn. (16), $\bm{\omega}\left(\bm{X},0\right)$ denotes the vorticity of a material point located in $\bm{X}$ in time 00 and $\bm{\omega}\left(\bm{\phi}\left(\bm{X}\right),t\right)$ is the vorticity of the same material point in time $t$. In order to establish the forward flow map Jacobian $\mathcal{F}$ that corresponds to the trajectory from position $\bm{X}$ and time 00 to position $\bm{x}$ and time $t$, we have to evolve $\mathcal{F}$ in reverse. The reverse evolution of $\mathcal{F}$ and $\mathcal{T}$ is governed by:
@@ 4.2. Revised Poisson Equation
It can be observed from the fact that $u_{x}^{i,j-1}$ is now dependent on $u_{y}^{i-1,j}$, that such a revised system of equations will be coupled across spatial dimensions, i.e., it can no longer be solved using a generic scalar Poisson solver by treating each dimension separately. For the general case in 3D, $u_{x}$ has six neighboring velocities, with two on each axis. The y and z-axis neighbors differ from the x-axis neighbors due to an intervening vorticity variable. If this variable is not within the fluid (e.g., on a solid boundary), we adjust the LHS by subtracting one from the coefficient of $u_{x}$, removing the corresponding velocity, and adding or subtracting unknowns from other axes within the fluid depending on their relative positions, as shown by an analogous 2D example in Figure 7. Then, we adjust the RHS by removing such intervening vorticities and moving the known velocities on the walls along the x-axis to the RHS. More details and algorithms are discussed in the Appendix Section A.
@@ 4.3. Coupled Poisson Solver
To solve this revised equation system efficiently and accurately, we develop a GPU-based matrix-free Multi-grid Preconditioned Conjugate Gradient (MGPCG) solver, whose implementation details are given in the Appendix Section A.1. Our solver is a geometric multigrid solver implemented in Taichi (Hu et al., 2019), with solid boundary conditions being enforced at coarser levels with coarsened solids. The restriction, prolongation, and smoothing are implemented following (McAdams et al., 2010). On one hand, as seen in Figure 9, our method is able to significantly improve the simulation realism with its physically correct treatment of vorticity-velocity coupling near the boundary. On the other hand, as shown in Table 6.1, by leveraging the matrix-free nature of our solver, we are able to conveniently, without incurring a hindering memory overhead, harness the parallel computation of modern GPUs to process multiple dimensions concurrently.
@@ 5.1. Discretization
8:     $\hat{\bm{\omega}}\leftarrow$ BFECC$(\bm{\omega}_{\text{init}},\bm{\psi},\mathcal{T},\bm{\phi},\mathcal{F})$;
@@
Input: $j$, $\mathcal{V}$, $\mathcal{S}$.
Output: $\bm{\psi}$, $\mathcal{T}$, $\bm{\phi}$, $\mathcal{F}$.
@@
3:March $\bm{\psi},\mathcal{F}$ with Algorithm 3, using $\bm{u}_{\text{mid}}$ and $-\Delta t_{j}$;
4:for $l$ in j-1 …0 do
5:     Fetch $\bm{u}_{\text{mid},l}$ and $\Delta t_{l}$ from the buffer $\mathcal{V}$ and $\mathcal{S}$;
6:     March $\bm{\psi},\mathcal{F}$ with Algorithm 3, using $\bm{u}_{\text{mid},l}$ and $-\Delta t_{l}$;
7:end for
8:March $\bm{\phi},\mathcal{T}$ with Algorithm 3, using $\bm{u}_{\text{mid}}$ and $\Delta t_{j}$.
@@ 5.2. Time Integration
Figure 11. A vortex pair successfully passes through a gap between two disks, demonstrating that by implicitly modeling the harmonic component, our result is consistent with Fluid Cohomology (Yin et al., 2023a), where the harmonic component is explicitly modeled. An experiment with a wider gap is also included in the video.
@@ 5.3. Implementational Details
We adopt the accumulation buffer from (Qu et al., 2019) for external forces and viscosity. In each step, we compute the current change with forward map $\bm{\phi}$ and add this change to to the buffer at its undeformed position.
Figure 12. Streamline of the velocity fields of the lid-driven cavity flow with the top boundary velocity being fixed at 1m/s horizontally right and the other three boundaries being static solid walls. Our method generates three vorticies (highlighted with red boxes), which is consistent with the ground truth (Ghia et al., 1982) under the same Re = 5000.
@@ 6.1. Solid Boundary Experiments and Divergence Test
Lid-driven cavity flow. Figure 12 shows an example of the 2D Lid-driven cavity flow where the velocity of the top boundary is fixed at 1m/s horizontally to the right, while the other three boundaries are solid walls. We tested our method with Re = 5000, and the result is consistent with the ground truth shown in (Ghia et al., 1982).
@@
Figure 16. Comparisons of 2D leapfrog vortices (top) and Taylor vortex (bottom). The columns from left to right depict ours, Covector Fluids (CF) (Nabizadeh et al., 2022), vorticity on flow map, Vortex-in-Cell, Semi-Lagrangian (SL) advection, Back and Forth Error Compensation and Correction (BFECC) respectively (SL and BFECC only refer to the advection techniques used, so these are still vortex methods. A normal velocity-pressure BFECC comparison is shown in Figure 20).
@@
Figure 19. A flapping stingray swimming at a constant speed, as simulated by our Eulerian vortex method. This figure shows the turbulent vorticity field generated by the stingray’s movement at different time frames.
@@ 6.2. Performance Analysis and Reinitialization Sensitivity
Flapping and static stingray. Figure 19 illustrates a stingray swimming steadily, creating vortices via flapping. Figure 17 (bottom) displays a static stingray with inflow under varying viscosities.

### v2

@@ Abstract.
We present an Eulerian vortex method based on the theory of flow maps to simulate the complex vortical motions of incompressible fluids. Central to our method is the novel incorporation of the flow-map transport equations for line elements, which, in combination with a bi-directional marching scheme for flow maps, enables the high-fidelity Eulerian advection of vorticity variables. The fundamental motivation is that, compared to impulse $\bm{m}$, which has been recently bridged with flow maps to encouraging results, vorticity $\bm{\omega}$ promises to be preferable for its numerical stability and physical interpretability. To realize the full potential of this novel formulation, we develop a new Poisson solving scheme for vorticity-to-velocity reconstruction that is both efficient and able to accurately handle the coupling near solid boundaries. We demonstrate the efficacy of our approach with a range of vortex simulation examples, including leapfrog vortices, vortex collisions, cavity flow, and the formation of complex vortical structures due to solid-fluid interactions.
@@
Figure 1. Turbulent flow generated from a delta wing with zero viscosity and an angle-of-attack of $20^{\circ}$. Two pictures on the right illustrate smoke (top) and vorticity (bottom) snapshots at different time steps. A more stable spiral vortex structure along the two leading edges can be generated when viscosity is considered, as shown in Figure 17 (top).
@@ 1. Introduction
Vortex methods have established themselves as a foundational computational method for incompressible flows with turbulence and complex vortical structures. In these methods, the Navier-Stokes equations written in the velocity form $\bm{u}$ are reformulated into their vorticity form, establishing vorticity
@@ 2. Related Work
Vortex methods rewrite the incompressible Navier-Stokes equation by treating vorticity as a gauge variable for velocity. By explicitly advecting the vorticity, this approach naturally preserves fluid circulation. In order to reduce the numerical dissipation during the advection, diverse vorticity representations have been proposed, ranging from particles (Selle et al., 2005; Cottet et al., 2000; Park and Kim, 2005; Zhang and Bridson, 2014; Angelidis, 2017), to filaments (Angelidis and Neyret, 2005; Weißmann and Pinkall, 2010; Ishida et al., 2022; Padilla et al., 2019), segments (Xiong et al., 2021), sheets (Brochu et al., 2012; Pfaff et al., 2012; Da et al., 2015), and Clebsch level sets (Chern et al., 2016, 2017; Yang et al., 2021; Xiong et al., 2022). For these Lagaragian vorticity representations, the velocity reconstruction from vorticity is often done with the Biot-Savart law, accelerated with the Fast Multipole Method (FMM) (Greengard and Rokhlin, 1987). Despite the enhanced ability to preserve the vortex structure, these vorticity representations are not as convenient to implement as the Eulerian vortex method when handling solid boundaries. Vortex methods with the vorticity-streamfunction formulation have been adopted by several researchers (Zhang et al., 2015; Elcott et al., 2007; Ando et al., 2015; Yin et al., 2023a). In spite of the ability to handle solid boundaries, these methods require solving a potential component of the velocity. Existing Eulerian vortex methods that use the velocity-vorticity formulation avoid the streamfunction but suffer from poor vorticity preservation and inaccurate solid boundary conditions. For example, Huang and Li (1997) only addressed 2D cases and Liu (2001) set the vorticity boundary condition simply as $\bm{\omega}_{b}=0$. Our method bypasses the use of streamfunction and potential without sacrificing the correctness in solid boundary handling, and pushes forward the state-of-the-art in vortex method simulations through the novel combination with flow map advection.
@@ 3.1. Flow map
Here, Eqn. (6) defines the forward evolution, and Eqn. (7) defines the time-reversed evolution. In Eqn (6), we have $\bm{\psi}\left(\bm{x}\right)$ and $\bm{x}$ specifying the start and end point of the flow map path at time 00 and $t$. Conversely, in Eqn (7), by examining the backward path from time $t$ to time 00, we can construct a mapping with $\bm{l}\left(\bm{X},0\right)$ denoting the line element located in $\bm{X}$ at time 00 and $\bm{l}\left(\bm{\phi}\left(\bm{X}\right),t\right)$ specifying its value at time $t$.
@@
(10) $\bm{s}(\bm{X},0)=\mathcal{F}^{T}\,\bm{s}\left(\bm{\phi}\left(\bm{X}\right),t\right).$
@@ 3.3. Vorticity on Flow Maps
Here, in Eqn. (15), $\bm{\omega}\left(\bm{x},t\right)$ denotes the vorticity of a material point located in $\bm{x}$ at time $t$ and $\bm{\omega}\left(\bm{\psi}\left(\bm{x}\right),0\right)$ is the vorticity of the same material point at time 00. Conversely, in Eqn. (16), $\bm{\omega}\left(\bm{X},0\right)$ denotes the vorticity of a material point located in $\bm{X}$ at time 00 and $\bm{\omega}\left(\bm{\phi}\left(\bm{X}\right),t\right)$ is the vorticity of the same material point at time $t$. In order to establish the forward flow map Jacobian $\mathcal{F}$ that corresponds to the trajectory from position $\bm{X}$ and time 00 to position $\bm{x}$ and time $t$, we have to evolve $\mathcal{F}$ in reverse. The reverse evolution of $\mathcal{F}$ and $\mathcal{T}$ is governed by:
@@ 4.2. Revised Poisson Equation
It can be observed from the fact that $u_{x}^{i,j-1}$ is now dependent on $u_{y}^{i-1,j}$, that such a revised system of equations will be coupled across spatial dimensions, i.e., it can no longer be solved using a generic scalar Poisson solver by treating each dimension separately. For the general case in 3D, $u_{x}$ has six neighboring velocities, with two on each axis. The y and z-axis neighbors differ from the x-axis neighbors due to an intervening vorticity variable. If this variable is not within the fluid (e.g., on a solid boundary), we adjust the LHS by subtracting one from the coefficient of $u_{x}$, removing the corresponding velocity, and adding or subtracting unknowns from other axes within the fluid depending on their relative positions, as shown by an analogous 2D example in Figure 7. Then, we adjust the RHS by removing such intervening vorticities and moving the known velocities on the walls to the RHS. More details and algorithms are discussed in the Appendix Section A.
@@ 4.3. Coupled Poisson Solver
To solve this revised equation system efficiently and accurately, we develop a GPU-based matrix-free Multi-Grid Preconditioned Conjugate Gradient (MGPCG) solver, whose implementation details are given in the supplementary material. Our solver is a geometric multi-grid solver implemented in Taichi (Hu et al., 2019), with solid boundary conditions being enforced at coarser levels with coarsened solids. The restriction, prolongation, and smoothing are implemented following (McAdams et al., 2010). On one hand, as seen in Figure 9 (top), our method is able to significantly improve the simulation realism with its physically correct treatment of vorticity-velocity coupling near the boundary. On the other hand, as shown in Table 6.1, by leveraging the matrix-free nature of our solver, we are able to conveniently, without incurring a hindering memory overhead, harness the parallel computation of modern GPUs to process multiple dimensions concurrently.
@@ 5.1. Discretization
8:     $\hat{\bm{\omega}}\leftarrow$ BFECC$(\bm{\omega}_{\text{init}},\bm{\psi},\mathcal{F},\bm{\phi},\mathcal{T})$;
@@
Input: $j$, $\mathcal{V}$, $\mathcal{S}$, $\bm{\phi}$, $\mathcal{T}$.
Output: $\bm{\psi}$, $\mathcal{F}$, $\bm{\phi}$, $\mathcal{T}$.
@@
3:for $l$ in j …0 do
4:     Fetch $\bm{u}_{\text{mid},l}$ and $\Delta t_{l}$ from the buffer $\mathcal{V}$ and $\mathcal{S}$;
5:     March $\bm{\psi},\mathcal{F}$ with Algorithm 3, using $\bm{u}_{\text{mid},l}$ and $-\Delta t_{l}$;
6:end for
7:March $\bm{\phi},\mathcal{T}$ with Algorithm 3, using $\bm{u}_{\text{mid}}$ and $\Delta t_{j}$.
@@ 5.2. Time Integration
Figure 11. A vortex pair successfully passes through a gap between two disks, demonstrating that by implicitly modeling the harmonic component, our result is consistent with Fluid Cohomology (Yin et al., 2023a), where the harmonic component is explicitly modeled.
@@ 5.3. Implementational Details
We adopt the accumulation buffer from (Qu et al., 2019) for external forces and viscosity. In each step, we compute the current vorticity change and add this change back to the buffer at its undeformed position with forward map $\bm{\phi}$.
@@ 6.1. Solid Boundary Experiments and Divergence Test
Lid-driven cavity flow. Figure 14 shows the streamline of the 2D Lid-driven cavity flow where the velocity of the top boundary is fixed at 1m/s horizontally to the right, while the other three boundaries are static solid walls. We tested our method with Re = 5000, it generates three vortices at the left top, left bottom and right bottom corner, which is consistent with the ground truth shown in (Ghia et al., 1982).
Figure 14. Cavity flow
@@
Figure 16. Comparisons of 2D leapfrog vortices (top) and Taylor vortex (bottom). The columns from left to right depict ours, Covector Fluids (CF) (Nabizadeh et al., 2022), the vorticity version of CF (traditional flow map advection), Vortex-in-Cell, Semi-Lagrangian (SL) advection, Back and Forth Error Compensation and Correction (BFECC) respectively (SL and BFECC only refer to the advection techniques used, so these are still vortex methods. A normal velocity-pressure BFECC comparison is shown in Figure 20).
@@
Figure 19. A stingray passively glides through water at a constant speed with zero viscosity, as simulated using our Eulerian vortex method. The figure illustrates the turbulent vorticity field generated at different time steps.
@@ 6.2. Performance Analysis and Reinitialization Sensitivity
Gliding stingray with viscosity. Figure 19 illustrates a stingray passively gliding, generating vortices. Figure 17 (bottom) displays the stingray with inflow under different viscosity conditions.